not positive if i have graphed correctly 5x+3y=15 need to make sure
Answer by Nate(3500)

Your Question:not positive if i have graphed correctly 5x+3y=15 need to make sure
--------------------------------------------------------------
My answer:
First you get the equation into y=mx+b (point-intercept form)
5x+3y=15-----original equation
3y=-5x+15-----subtract -5x
y=(-5/3)x+5-----divide everything by three to get y by itself
In y=mx+b, m=slope and b=y-intercept.
To graph, first graph 5 units above the origin.
Then, either draw another dot 5 units up and 3 units to the left of the point (0,5) or 5 units down and 3 units to the right(this is using your slope of -5/3.)
Connect the dots and that is your graph!